An electron is located on the x axis at x0 = -1.33 × 10^-6 m. Find the magnitude and direction of the electric field at x = 3.63 × 10^-6 m on the x axis due to this electron. Magnitude:
Answer #1

E=kq/r2

q=- 1.6*10-19C , r= (3.63*10-6 –(-1.33*10-6))=

E= (9*109* -1.6*10-19)/(4.99*10-6)2 = - 57.83 N/C

Direction of E is towards to center of the electron.